OpenStudy (anonymous):

0.05x+0.25y=11 0.15x+0.05y=12 order pair again I have five I could get out of 20 please help

OpenStudy (anonymous):

Didn't believe me?

OpenStudy (anonymous):

you said that you thought that was it not sure is what I got sorry....

OpenStudy (anonymous):

No no not that, I gave you the approximate answer

OpenStudy (anonymous):

There will be several decimal places and all

OpenStudy (anonymous):

I ignored them

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

correct thanks and sorry again. next

OpenStudy (anonymous):

multiply the 2nd equation by 5 0.05x+0.25y=11 0.75x+0.25y=60 subtract eq2 from eq1 -0.7x=-49 x=-49/-0.7 x=70 now replace 70 in x in any eq and solve y 0.05*70+0.25y=11 y=30

OpenStudy (anonymous):

did zat help u?